OpenStudy (anonymous):

The running time for movie A is 36 minutes more than twice the running time for a movie B. If the running time for movie A is subtracted from triple the running time for movie B, the result is 58 minutes. What is my answer? Running time movie A is ___ mins and running time for movie B is ___ mins.

OpenStudy (anonymous):

Can I help???????????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I just need to figure this question

OpenStudy (anonymous):

whait i will come back in a second

OpenStudy (anonymous):

i am back

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

what do you think the answere is

OpenStudy (anonymous):

@KendyM25

OpenStudy (anonymous):

Well I haven't bit problem trying to figure it out its algebra question that complete on lab quiz

OpenStudy (danjs):

Translate each sentence into an equation.. use A and B for the running time variable for each is - =

OpenStudy (anonymous):

A is 36 minutes more than twice the running time for a movie B right here we have one clue

OpenStudy (danjs):

The running time for movie A --- A is --- = 36 minutes more than --- + 36 twice the running time for a movie B --- 2*B ------------------------------------------

OpenStudy (anonymous):

If the running time for movie A is subtracted from triple the running time for movie B, the result is 58 minutes here is the second

OpenStudy (danjs):

the running time for movie A is subtracted from triple the running time for movie B --- B-A the result is --- = 58 minutes ---- 58 ,

OpenStudy (anonymous):

Yea this math is algebraic expression of problem values

OpenStudy (anonymous):

Idk if anyone is familiar with consecutive integers problems cause that's what this is.

OpenStudy (anonymous):

did you find out the answere

OpenStudy (danjs):

"The running time for movie A is 36 minutes more than twice the running time for a movie B" A = 2*B + 36 " If the running time for movie A is subtracted from triple the running time for movie B, the result is 58 minutes" B - A = 58

OpenStudy (danjs):

They gave you two equations that include the running times for A and B. You can solve that system of two equations for the two variables A and B A = 2*B + 36 B - A = 58

OpenStudy (danjs):

Graphically, those are two lines in the coordinate plane. Each point is, (A,B). The solution is the intersection of both solution sets, where the two lines cross, at the point (A,B), A= and B= some value.